OpenStudy (pig_cat223):

PLZ HELP!!! im really bad at this... solve this system of linear equations. seperate x- and y-values with a comma. 2x=-24-4y -5x=36=4y

OpenStudy (marcelie):

same process :)

OpenStudy (pig_cat223):

okay... with out using my problem can you explain the substitution to me?

OpenStudy (marcelie):

yes , im going to pick number one equation so my goal is to solve for y

OpenStudy (pig_cat223):

okay?...

OpenStudy (marcelie):

so after u get ur y ur going to substituted for your second equation to find ur x

OpenStudy (marcelie):

2nd equation whats the sign ? bc thres two =

OpenStudy (pig_cat223):

the sign?

OpenStudy (marcelie):

yes i think u made typo

OpenStudy (marcelie):

-5x=36=4y?

OpenStudy (pig_cat223):

oh yeah i did sorry lol.... its +

OpenStudy (marcelie):

which one lol

OpenStudy (marcelie):

-5x+36=4y?

OpenStudy (pig_cat223):

the second one.... its suppose to be... -5x=36+4y

OpenStudy (marcelie):

okay.. so ur first equation will look like y= -1/2x-6

OpenStudy (marcelie):

you get that part?

OpenStudy (marcelie):